# Predictability of viral load kinetics in the early phases of SARS-CoV-2 through a model-based approach

### v1

@@ 3.4 Inference of viral load dynamics based on uncertain data
Step 2 – Inclusion of the uncertainty. We now go further in the analysis by including the presence of uncertainty, by looking at the full model (2). The aim is then to perform an uncertainty quantification on the shape of $\beta(\tau,\mathbf{z}_{j})$, with respect to the randomness $(\mathbf{z}_{j})_{j=1}^{M}$ that may appear when estimating $\alpha_{1}(\mathbf{z}_{j})$, the day of first contact. For this, we perturb the optimal value $\alpha_{1}$ computed in the previous step by a centered uniform random distribution $\mathcal{U}([-2,2])$, i.e. $\alpha_{1}(\mathbf{z}_{j})=\alpha_{1}+\mathbf{z}_{j}$ with $\mathbf{z}_{j}\sim\mathcal{U}([-2,2])$, and we assess the impact of this addition on the parameters $\alpha_{2}(\mathbf{z}_{j})$ and $\alpha_{3}(\mathbf{z}_{j})$ that define the shape of the $\beta(\tau,\mathbf{z}_{j})$. The (randomized) optimal shape is thus determined by fitting model (2) to reproduce the data on infected and recovered cases, through a minimization problem similar to (5), but with parameters $\alpha_{1}(\mathbf{z}_{j})$ and $\alpha_{2}(\mathbf{z}_{j})$ instead of $\alpha_{1}$ and $\bar{\beta}$, while of course the latter and $\bar{\gamma}$ are those obtained from the previous step. This procedure is repeated for the whole sample $\mathcal{C}_{z}$, and we can finally infer the optimal shape of the infectiousness function with UQ, to be the expected value $\mathbb{E}_{z}(\beta(\tau,\mathbf{z}))$ with respect to $\mathbf{z}$ and based on the probability density $p(\mathbf{z})=\mathcal{U}([-2,2])(\mathbf{z})$.
@@
Figure 3: Data- and model-driven inference of the optimal shape of the viral load in the province of Pavia (Italy), over two different epidemic waves: the one occurred at the beginning of the pandemic in 2020, with a focus on the two weeks (February 24–March 9) before the first lockdown was established (top), and the other occurred in the fall of the same year, with a focus on the two weeks (October 7–22) preceding the imposition of a national curfew (bottom). The figures on the left provide the optimal viral load shapes (red curves) determined for these two periods, starting from the data collected by Fondazione IRCCS Policlinico San Matteo. The shaded areas show how the infectiousness function is altered when a uniform noise $\mathcal{U}([-2,2])$ modifies the day of first contact. The figures on the right present the evolution of infected from SARS-CoV-2 in the province of Pavia over the considered windows of time. We compare computed (in red) and reported (in black) number of cases.
The outcomes of our two-step procedure are reported in Figure 3 for the case of Pavia and in Figure 4 for the case of Italy. In the left figures we plot in red the curves providing the mean best shape of the infectiousness function over the three epidemic waves that we have considered. The shaded regions show the $80\%$ and $95\%$ percentiles of the different shapes, based on the sample $\mathcal{C}_{z}$. We then perform numerical simulations to assess the viability of the model (2) to predict the evolution of infected individuals during these subsequent phases of the pandemic. We take the infectiousness function to be the computed mean best shapes over the three periods and we choose $\gamma(\tau,\mathbf{z})$ constant and equal to 1 for simplicity. In particular, the duration of the infection is solely determined by $\bar{\gamma}$ and the contact rate is given by the optimal ones $\bar{\beta}$ determined from Step 1 and reported in Table 3. We provide in Figures 3 and 4, on the right, the comparisons between the reported number of infected in Pavia and in Italy over the three epidemic waves, and the corresponding predictions of the model (2), with the aforementioned choices. We notice that with the progression of the pandemic, and specifically over the third wave in March 2021, the model loses some accuracy since it is less appropriate to describe slower growths of the epidemic, stemming from the numerous containment measures introduced by the Italian government in the course of 2021 to reduce contacts among the population. This suggests in particular that in order to capture such intricate effects and minimize the error made by the model, one should consider a variable contact rate $\bar{\beta}(t)$ that continuously evolves over the epidemic time $t$.
@@
Figure 4: Investigations on the SARS-CoV-2 pandemic in Italy, during three distinct epidemic waves. Top row: analysis of the epidemic wave that started in February–March 2020. Optimal shape $\mathbb{E}_{z}(\beta(\tau,\mathbf{z}))$ of the infectiousness function (left) and model predictions (right) for the cases of infection registered during the period February 24–March 9. Middle row: analysis of the new epidemic outburst that took place in October 2020. Optimal shape of the viral load (left) and model predictions (right) for the cases of infection registered during the period October 7–22. Bottom row: analysis of the epidemic wave occurred from March 2021. Optimal shape of the viral load (left) and model predictions (right) for the cases of infection registered during the period February 24–March 10.
@@
Extending the model predictions over the whole timeframe. We now perform a calibration of our model to cover the whole timespan of observation, connecting the epidemic waves between them and providing full predictions for the evolution of infected. We initially observe that as the epidemiological parameter $\bar{\beta}$ models the contact rate between individuals, its value is highly influenced by the restrictions imposed by a government. This fact needs to be taken into account when extending our model’s predictions from one epidemic wave to the other, since various containment measures (we mention in particular the lockdown imposed between March and May 2020) have been put in place by the Italian government over time to slow down the epidemic spreading. Since we do not need to evaluate the shape of the viral load in these connecting timeframes (the previous two-step analysis is relevant only when the virus is able to spread free among the population, allowing for a proper calibration of the epidemiological parameters), in order to connect the observations we can rely on a simpler model that allows to speed up the computational time. Specifically, we make use of system (2) with the choices $\beta(\tau,\mathbf{z}),\ \gamma(\tau,\mathbf{z})\equiv 1$. From the above considerations, we account for the variable contact restrictions by assuming that the contact rate varies over the epidemic time as $\bar{\beta}=\bar{\beta}(t)$. Then, we determine the value $\bar{\beta}(t)$ by solving an optimization problem for a sequence of time steps $t_{i}$ (days) over a moving time window of three days (namely, averaging the fitting over three days), varying in any timespan $t\in[t_{0}+1,t_{f}]$ where $t_{0}$ and $t_{f}$ correspond respectively to the end of an epidemic wave and the beginning of the next one. This strategy is needed in order to correctly assess the evolution of the contact rate depending on the containment measures that are in place. More precisely, we solve the following constrained least-square problem
@@
and the recovery rate $\bar{\gamma}$ is always kept fixed to the value $1/14$. We complete our study by pushing the model predictions up to the last observation times of our experimental data on the epidemic, that are January 18 2021 for the case of the province of Pavia and August 4 2021 for the case of Italy. Using the above procedure, we connect the observations by calibrating $\bar{\beta}(t)$ with the minimization strategy (6). Using all these optimized parameters, we can finally run our model over the whole timeframe of observation. We depict the resulting predictions for the evolution of infected individuals in the province of Pavia, in Figure 5 and Figure 6 respectively, on the right. The confirmed cases are plotted in black, while the model approximations are shown in red.
@@
Figure 5: Left: comparison between viral loads from the first (in orange, March 2020) and the second (in blue, October 2020) epidemic waves, in the province of Pavia. In order for the SIR model (2) to best fit the data on confirmed cases, the peak of the infection should be reached about 6 to 8 days after the first contact with an infected subject. Right: evolution of infected from SARS-CoV-2 in Pavia from February 24 2020 to January 18 2021. Comparison between computed (in red) and reported (in black) number of cases.
@@
At last, in Figures 5 and 6 on the left, we provide a direct comparison between the optimal viral loads from different epidemic waves, obtained via the previous two-step analysis, together with their $80\%$ prediction bounds. These are aligned by their respective peak of infectiousness (which corresponds to day 0 on the graph). One can observe that the infection’s peak progressively decreases, while the tails of the viral load tend to grow. This suggests that the virus becomes weaker but more effective over a longer period of time, and individuals stay infective for more days on average. Moreover, the decrease in curve’s growth in the beginning of the infection could suggest a slower progression of the symptoms and explain the greater spreading of the epidemic with each new wave.
@@
Figure 6: Left: comparison between viral loads from three distinct epidemic waves in Italy, happened respectively in March 2020 (orange), in October 2020 (blue) and in March 2021 (green). In order for the SIR model (2) to best fit the data on confirmed cases, the peak of the infection should be reached about 5 to 8 days after the first contact with an infected subject. Right: evolution of infected from SARS-CoV-2 in Italy from February 24 2020 to August 4 2021. Comparison between computed (in red) and reported (in black) number of cases.
4 Discussion
@@
Understanding the structure and evolution of VL kinetics is crucial in the context of public health, when setting up quarantine and isolation measures of positive subjects. In particular, one of the critical issues emerged since the onset of the recent SARS-CoV-2 pandemic has been the need of criteria to assess the infectivity of patients and decide their releasing from prolonged quarantine, safely admitting them back to work duties and social activities. It was demonstrated that high Ct values (low VL), greater than 30-35 Ct, were associated with a reduced infectivity of the patients [29, 39, 35]. Nevertheless, existing studies indicate that the peak viral load is linked closely with the onset of symptoms making the available datasets partial and of difficult interpretation. In this work, we have established a pipeline to assess the evolution of in-host VL from respiratory samples of patients. To this end, we combined the use of real data and model-based methods. Furthermore, the provided method is capable of tracing the whole VL kinetics. In particular, our model helps to provide a data-oriented understanding of the non-infectious status of people that were under quarantine with high Ct.
@@
Samples included in this study were collected over two periods of time (November–December 2020 and January–May 2021), that include the initial circulation of VOCs and before the extensive introduction of vaccination. We have first determined the shape of the average VL kinetics hinted by these data (recovering Gamma-type distributions) and we have successively introduced a suitable compartmental model of SIR-type, using a local incidence rate function that varies with the age of infection, to account for the evolution of the epidemic alongside that of the virus. The system of equations belongs to a large class of epidemiological models detailing the transmission dynamics through the inclusion of additional factors such as external influence, age structure, variable contact dynamics and mobility of agents [1, 2, 9, 18, 27]. In this work, we exploit the information coming from real data on molecular tests and on the reported number of infected individuals to deduce the optimal shape of the infectivity function, based on a combined data-driven and model-driven approach. To take into account the extreme variability of VL dynamics, depending on disease severity and patients’ features as well [31, 44], we also make use of UQ methods [24, 34, 41] adapted for the study of epidemic systems [1, 10, 13, 38], to take into account the uncertainties inherent to the data-assimilation processes of VL kinetics. The model allows us to infer a plausible evolution of the SARS-CoV-2 infectiousness over three distinct periods of the pandemic (February–March 2020, October 2020 and February–March 2021), in agreement with existing works from the medical literature [11, 37]. In particular, we observed that the peak of the infection function decreases over time while its tails tend to grow, suggesting that the virus becomes progressively weaker but more effective, since individuals remain infective for more days on average. This change in shape strongly manifests during the third observation period (February–March 2021) when the Alpha variant was actively circulating in Italy, providing an explanation for the greater spreading of the contagion at the time.
@@ Acknowledgements
A.B. and M.Z. acknowledge the support from GNFM of INdAM (National Institute of High Mathematics) and from the Italian Ministry of University and Research (MUR) through the PRIN 2020 project (No. 2020JLWP23) “Integrated Mathematical Approaches to Socio–Epidemiological Dynamics”. A.B. also acknowledges the support from the European Union’s Horizon Europe research and innovation programme, under the Marie Skłodowska-Curie grant agreement No. 101110920 “MesoCroMo – A Mesoscopic approach to Cross-diffusion Modelling in population dynamics”. M.Z. wishes to acknowledge partial support by ICSC - Centro Nazionale di Ricerca in High Performance Computing, Big Data and Quantum Computing, funded by European Union - NextGenerationEU.

### v2

@@ 3 Inference of viral load dynamics
Step 2 – Inclusion of the uncertainty. We now go further in the analysis by including the presence of uncertainty and by looking at the full model (3). The aim is to perform an uncertainty quantification on the shape of $\beta(\tau,\mathbf{z}_{j})$ with respect to the randomness $(\mathbf{z}_{j})_{j=1}^{M}$ that may appear when estimating $\alpha_{1}(\mathbf{z}_{j})$, the day of first contact. For this, we perturb the optimal value $\alpha_{1}$ computed in the previous step by a centered uniform random distribution $\mathcal{U}([-2,2])$, i.e. $\alpha_{1}(\mathbf{z}_{j})=\alpha_{1}+\mathbf{z}_{j}$ with $\mathbf{z}_{j}\sim\mathcal{U}([-2,2])$, and we assess the impact of this addition on the parameters $\alpha_{2}(\mathbf{z}_{j})$ and $\alpha_{3}(\mathbf{z}_{j})$ that define the shape of $\beta(\tau,\mathbf{z}_{j})$. The (randomized) optimal shape is thus determined by fitting the model (3) to reproduce the data on infected and recovered cases, through a minimization problem similar to (6), but with parameters $\alpha_{1}(\mathbf{z}_{j})$ and $\alpha_{2}(\mathbf{z}_{j})$ instead of $\alpha_{1}$ and $\bar{\beta}$, while of course the latter and $\bar{\gamma}$ are those obtained from the previous step. This procedure is repeated for the whole sample $\mathcal{C}_{z}$ and we can finally infer the optimal shape of the infectiousness function with UQ to be the expected value $\mathbb{E}_{z}(\beta(\tau,\mathbf{z}))$ with respect to $\mathbf{z}$, based on the probability density $p(\mathbf{z})=\mathcal{U}([-2,2])(\mathbf{z})$.
@@
Figure 3: Data- and model-driven inference of the optimal shape of the infection rate in the province of Pavia (Italy) over two different epidemic waves: the one occurred at the beginning of the pandemic in 2020, with a focus on the two weeks (February 24–March 9) preceding the first lockdown (top), and the other occurred in the fall of the same year, with a focus on the two weeks (October 7–22) preceding the imposition of a national curfew (bottom). The figures on the left provide the optimal shapes $\mathbb{E}_{z}(\beta(\tau,\mathbf{z}))$ of the infectiousness function (red curves) determined for these two periods, starting from the data collected by Fondazione IRCCS Policlinico San Matteo to infer an initial guess of the infection rate through the function (2) rescaled by the factor $\alpha_{0}=10^{7}$. The shaded areas show how the infectiousness function is altered when a uniform noise $\mathcal{U}([-2,2])$ modifies the day of first contact. The figures on the right present the evolution of infected individuals from SARS-CoV-2 in the province of Pavia over the considered windows of time. We compare computed (in red) and reported (in black) number of cases.
@@
Figure 4: Investigations on the SARS-CoV-2 pandemic in Italy during three distinct epidemic waves, starting from the data collected by Fondazione IRCCS Policlinico San Matteo to infer an initial guess of the infection rate through the function (2). Top row: analysis of the epidemic wave that started in February–March 2020. Optimal shape $\mathbb{E}_{z}(\beta(\tau,\mathbf{z}))$ of the infectiousness function (left) and model approximation (right) for the cases of infection registered during the period February 24–March 9 2020. Middle row: analysis of the new epidemic outburst that took place in October 2020. Optimal shape of the infection rate (left) and model approximation (right) for the cases of infection registered during the period October 7–22 2020. Bottom row: analysis of the epidemic wave occurred from March 2021.
The outcomes of our two-step procedure are reported in Figure 3 for the case of Pavia and in Figure 4 for the case of Italy. In the left figures we plot in red the curves providing the mean best shape of the infectiousness function over the three epidemic waves that we have considered. The shaded regions show the $80\%$ and $95\%$ percentiles of the different shapes, based on the sample $\mathcal{C}_{z}$. We then perform numerical simulations to assess the viability of model (3) to predict the evolution of infected individuals during these subsequent phases of the pandemic. We take the infectiousness function to be the computed mean best shapes over the three periods and we choose $\gamma(\tau,\mathbf{z})$ constant and equal to 1 for simplicity. In particular, the duration of the infection is solely determined by $\bar{\gamma}$ and the contact rate is given by the optimal ones $\bar{\beta}$ determined from Step 1 and reported in Table 3. We show in Figures 3 and 4, on the right, the comparisons between the reported number of infected individuals in Pavia and in Italy over the three epidemic waves and the corresponding approximations provided by model (3), with the aforementioned choices of the parameters. We notice that with the progression of the pandemic, specifically over the third wave of February–March 2021, the model loses some accuracy since it is less appropriate to describe slower growths of the epidemic, which stem from the numerous containment measures introduced by the Italian government in the course of 2021 to reduce contacts among the population. This suggests in particular that, in order to capture such intricate effects and minimize the error made by the model, one should consider a variable contact rate $\bar{\beta}(t)$ that continuously evolves over the epidemic time $t$.
@@
Reconstructing the experimental data over the whole timeframe. We now perform a calibration of our model to cover the whole timespan of observation, connecting the epidemic waves between them and providing a full reconstruction of the experimental data on the evolution of infected individuals. We initially observe that, as the epidemiological parameter $\bar{\beta}$ models the contact rate between individuals, its value is highly influenced by the restrictions imposed by a government. This fact needs to be taken into account when extending our model’s predictions from one epidemic wave to the other, since various containment measures (we mention in particular the lockdown imposed between March and May 2020) have been put in place by the Italian government over time to slow down the epidemic spreading. Since we do not need to evaluate the shape of the infection rate in these connecting timeframes (the previous two-step analysis is relevant only when the virus is able to spread free among the population, allowing for a proper calibration of the epidemiological parameters), in order to connect the observations we can rely on a simpler model that allows to speed up the computational time. Specifically, we make use of system (3) with the choices $\beta(\tau,\mathbf{z}),\gamma(\tau,\mathbf{z})\equiv 1$. From the above considerations, we account for the variable contact restrictions by assuming that the contact rate varies over the epidemic time as $\bar{\beta}=\bar{\beta}(t)$. Then, we determine the value $\bar{\beta}(t)$ by solving an optimization problem for a sequence of time steps $t_{i}$ (days) over a moving time window of three days (namely, averaging the fitting over three days), varying in any timespan $t\in[t_{0}+1,t_{f}]$ where $t_{0}$ and $t_{f}$ correspond respectively to the end of an epidemic wave and the beginning of the next one. This strategy is needed in order to correctly assess the evolution of the contact rate depending on the containment measures that were in place. More precisely, we solve the following constrained least-square problem
@@
where the recovery rate $\bar{\gamma}$ is still kept fixed to the value $1/14$. We complete our study by using the model to reconstruct the experimental data on the epidemic up to the last observation times, being January 18 2021 for the case of the province of Pavia and August 4 2021 for the case of Italy. By means of the above procedure we connect the observations by calibrating $\bar{\beta}(t)$ with the minimization strategy (7). Using all these optimized parameters we can finally run our model over the whole timeframe of observation. We depict the resulting approximations for the evolution of infected individuals in the province of Pavia and in Italy, in Figure 5 and Figure 6 respectively, on the right. The evolution of the confirmed cases is plotted in black while the model approximations are shown in red.
@@
Figure 5: Left: comparison between infection rates from the first (in orange, March 2020) and the second (in blue, October 2020) epidemic waves in the province of Pavia. In order for the model (3) to best fit the data on confirmed cases, the peak of the infection should be reached about 6 to 8 days after the first contact with an infected subject. Right: evolution of infected individuals from SARS-CoV-2 in Pavia from February 24 2020 to January 18 2021. Comparison between computed (in red) and reported (in black) number of cases.
@@
At last, in Figures 5 and 6 on the left, we provide a direct comparison between the optimal infection rates from different epidemic waves, obtained via the previous two-step analysis, together with their $80\%$ prediction bounds. These are aligned by their respective peak of infectiousness (which corresponds to day 00 on the graph). One can observe that the infection’s peak progressively decreases, while the tails of the infection rate tend to grow. This suggests that the virus becomes weaker but more effective over a longer period of time and individuals stay infective for more days on average. Moreover, the decrease in curve’s growth in the beginning of the infection could suggest a slower progression of the symptoms and explain the greater spreading of the epidemic with each new wave.
@@
Figure 6: Left: comparison between infection rates from three distinct epidemic waves in Italy happened respectively in March 2020 (orange), in October 2020 (blue) and in March 2021 (green). In order for the model (3) to best fit the data on confirmed cases, the peak of the infection should be reached about 5 to 8 days after the first contact with an infected subject. Right: evolution of infected from SARS-CoV-2 in Italy from February 24 2020 to August 4 2021. Comparison between computed (in red) and reported (in black) number of cases.
4 Conclusion
@@
Understanding the structure and evolution of VL dynamics is crucial in the context of public health when setting up quarantine and isolation measures of positive subjects. In particular, one of the critical issues emerged since the onset of the recent SARS-CoV-2 pandemic has been the need of criteria to assess the infectivity of patients and decide their releasing from prolonged quarantine, safely admitting them back to work duties and social activities. It was demonstrated that high Ct values (low VL), greater than 30-35 Ct, were associated with a reduced infectivity of the patients [32, 46, 42]. Nevertheless, existing studies indicate that the peak viral load is closely linked with the onset of symptoms, making the available datasets partial and of difficult interpretation. In this work, we have established a pipeline to assess the evolution of in-host VL from respiratory samples of patients. To this end, we combined the use of real data and model-based methods. Furthermore, the proposed approach is capable of tracing the whole VL dynamics. In particular, our model helps to provide a data-oriented understanding of the non-infectious status of people that were under quarantine with high Ct.
@@
Samples included in this study were collected over two periods of time (November–December 2020 and January–May 2021) that include the initial circulation of VOCs and before the extensive introduction of vaccination. We have first determined the shape of the average VL dynamics hinted by these data (recovering Gamma-type distributions) and we have successively introduced a suitable compartmental model of SIR-type, using a local incidence rate function that varies with the age of infection, to account for the evolution of the epidemic alongside that of the virus. The system of equations belongs to a large class of epidemiological models detailing the transmission dynamics through the inclusion of additional factors such as external influence, age structure, variable contact dynamics and mobility of agents [1, 2, 11, 20, 30]. In this work, we exploited the information coming from real data on molecular tests and on the reported number of infected individuals to deduce the optimal shape of the infectivity function, based on a combined data-driven and model-driven approach. To take into account the extreme variability of VL dynamics, depending on disease severity and patients’ features as well [34, 51], we also made use of UQ methods [27, 39, 48] adapted for the study of epidemic systems [1, 12, 15, 45], to take into account the uncertainties inherent to the data-assimilation processes of VL kinetics. The model allowed us to infer a plausible evolution of the SARS-CoV-2 infectiousness over three distinct periods of the pandemic (February–March 2020, October 2020 and February–March 2021), in agreement with existing works from the medical literature [13, 44]. In particular, we observed that the peak of the infection function decreases over time while its tails tend to grow, suggesting that the virus becomes progressively weaker but more effective, since individuals remain infective for more days on average. This change in shape strongly manifests during the third observation period (February–March 2021) when the Alpha variant was actively circulating in Italy, providing an explanation for the greater spreading of the contagion at the time.
@@ Acknowledgements
A.B. and M.Z. acknowledge the support from GNFM of INdAM (National Institute of High Mathematics) and from the Italian Ministry of University and Research (MUR) through the PRIN 2020 project No. 2020JLWP23 (Integrated Mathematical Approaches to Socio-Epidemiological Dynamics). A.B. also acknowledges the support from the European Union’s Horizon Europe research and innovation programme, under the Marie Skłodowska-Curie grant agreement No. 101110920 (MesoCroMo - A Mesoscopic approach to Cross-diffusion Modelling in population dynamics). M.Z. wishes to acknowledge partial support by ICSC - Centro Nazionale di Ricerca in High Performance Computing, Big Data and Quantum Computing, funded by European Union - NextGenerationEU.